OpenStudy (anonymous):

In the exponential growth function y = 235(1.15)x, what is the: 1. growth factor? 2. percent increase?

OpenStudy (anonymous):

\[increase = 235(1.15)x - x= (235(1.15) - 1 )x= 270.25 x\]\[\%increase = \frac {270.25 x}{x} \times 100= 270.25 \%\]

OpenStudy (anonymous):

wait, is that \[(1.15)^x\]?

OpenStudy (anonymous):

oops, sorry I wrote it wrong. It is ^x

OpenStudy (anonymous):

I think I got the answer.